If the odds against an event happening are 81/55, what is the probability of the event happening?
Answer by stanbon(75887)

If the odds against an event happening are 81/55, what is the probability of the event happening?
-----
Note: odds against = P(against)/P(for)
---
P(for) = 55/(55+81) = 55/136 = 0.404..
-------
Cheers,
Stan H.
-----